A group hiked to a state park at 4mph. They took a truck back to town at 20mph. The round trip took 18 hours. How far is it from town to the park?
Found 3 solutions by richwmiller, checkley77, stanbon:
Answer by richwmiller(17219)

The distance to and from town is the same.
R*T=d
4t=20(18-t)
t=15
4*15=60 miles
15 hours to hike and 3 hours to ride the 60 miles

Answer by checkley77(12844)

D=RT
D=4X THE HICKING TRIP.
D=20(18-X) THE TRUCK TRIP.
4X=20(18-X)
4X=360-20X
4X+20X=360
24X=360
X=360/24
X=15 HOURS FOR THE HICKING TRIP.
18-15=3 HOURS FOR THE TRUCK TRIP.
PROOF:
4*15=20*3
60=60

Answer by stanbon(75887)

A group hiked to a state park at 4mph. They took a truck back to town at 20mph. The round trip took 18 hours. How far is it from town to the park?
------------------------------
Hike DATA:
rate = 4 mph ; distance = x miles ; time = x/4 hrs.
------------------
Truck DATA:
rate = 20 mph ; distance = x miles ; time = x/20 hrs.
------------------
Equation:
time + time = 18 hrs
x/4 + x/20 = 18
Multiply thru by 20 to get:
5x + x = 20*18
6x = 20*18
x = 60 miles (distance from town to the park)
============================
Cheers,
Stan H.